At 580nm potassium ferrocyanide has a molar absorptivity of 7000L(cm*mol) .What is the absorption of a 4,67 micromolar solution. I want to know when 4.67 absorption of the Beer's law how 580nm absorptivity of 7000 L/CM*1 MOL is calculated I don't remember I know we have the absorption already which is 4.67
Answer #1

According to Beer's law:

A = \epsiloncl

Where 'A' is the absorbance of the solution

'\epsilon' is the molar absorptivity of potassium ferrocyanide

'c' is the concentration of the solution

'l' is the path length = 1 cm (In general)

i.e. A = 7000 L/mol.cm * 4.67*10-6 mol/L * 1 cm (Note: 1 uM = 10-6 M, M = mol/L)

Therefore, the absorbance of the solution (A) = 0.03269 ~ 0.033